two salesman sold $75.80 worth of books.
one salesman sold $8.50 more than the other
How much did each salesman sell?

Answer by fcabanski(1391)

Call them salesman A and B. From the given facts:


A + B = 75.80 and A = B+8.50


B+8.50 + B = 75.80


2B = 67.30


B = 33.65


A = 33.65 + 8.50 = $42.15

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)